OpenStudy (anonymous):

(61,20) (72,34) Write the equation of the line passing through those two points using the point-slope formula y - y1 = m(x - x1). Show all of your work. Remember to find the slope of the line first.

OpenStudy (anonymous):

\[\huge slope = \frac{y_2-y_1}{x_2-x_1}=\frac{34-20}{72-61}=\frac{14}{11}\] standard equation y = mx + b substitute slope in y = (14/11) x + b substitute a point (20, 61) in 61 = (14/11)20 + b \[b =61-\frac{280}{11}=\frac{391}{11}\] y = (14/11)x + 391/11

OpenStudy (anonymous):

thank you